# Asymptotics for partitions over the Fibonacci numbers and related sequences

### v2

@@ 4. Partitions over general linear recurrences with a dominant root
We adopt the terminology of Proposition 8 for the rest of this section along with the definitions of the real numbers $\lambda=\lambda_{1},\lambda_{2},\ldots,\lambda_{r}$ satisfying
@@
where $C_{1}:=\sum_{k\geqslant 1}\sum_{\boldsymbol{k}\neq\boldsymbol{0}}\frac{1}{k_{1}}{k_{1}\choose k_{2}}\cdots{k_{r-2}\choose k_{r-1}}\frac{(\lambda_{r}\beta_{r}^{k})^{k_{r-1}}}{(\lambda\beta^{k})^{z+k_{1}}}\left(\prod_{j=2}^{r-1}(\lambda_{j}\beta_{j}^{k})^{k_{j+1}-k_{j}}\right)$.∎
@@
$s^{-z}\Gamma(z)\zeta(1+z)=\frac{1}{z^{2}}+\left(2\gamma-\log s\right)\frac{1}{z}+\left(\frac{1}{2}\left(\gamma^{2}+\frac{\pi^{2}}{6}\right)+\gamma^{2}-\gamma_{1}-2\gamma\log s+\frac{(\log s)^{2}}{2}\right)+O(z),$
@@
$\qquad-\left(\frac{\log\lambda}{\log\beta}+\frac{1}{2}\right)(2\gamma-\log s)+\left(\frac{(\log\lambda)^{2}}{2\log\beta}+\frac{\log\lambda}{2}+\frac{\log\beta}{12}+C_{1}\right)$
$=\frac{(\log s)^{2}}{2\log\beta}+\left(\frac{\log\lambda}{\log\beta}+\frac{1}{2}-\frac{2\gamma}{\log\beta}\right)\log s+C_{2},$
where $C_{2}:=\frac{1}{\log\beta}\left(\frac{1}{2}\left(\gamma^{2}+\frac{\pi^{2}}{6}\right)+\gamma^{2}-\gamma_{1}\right)-\left(\frac{\log\lambda}{\log\beta}+\frac{1}{2}\right)2\gamma+\left(\frac{(\log\lambda)^{2}}{2\log\beta}+\frac{\log\lambda}{2}+\frac{\log\beta}{12}+C_{1}\right).$ The contributions coming from the rest of the poles are as discussed before the statement of Lemma 4. ∎
On inspection, one notices that the dominant asymptotic terms of $\log F_{P}(e^{-s})$ are precisely the dominant asymptotic terms of $\log F_{2}(e^{-s})$, the function related to the Fibonacci partitions, after substituting $\beta$ for $\varphi$. Of course, this is not so surprising, as the Fibonacci numbers $F_{k}$ are just a special case of the more general sequence $P_{k}$. The property of note, here, is regarding the associated saddle point. Since the leading order behaviour is the same, the saddle points satisfy the same leading order asymptotics. In particular, Proposition 6 and Corollary 1 hold for the saddle point $\alpha_{F_{P}}$ related to non-distinct partitions of $n$ over $P_{k}$, and so also then, do (20) and (21). We use these results, as well as this notation, below.
@@
By Proposition 9, Proposition 6 and Corollary 1, the small-$\alpha_{F_{P}}$ asymptotics, or, equivalently, the large-$n$ asymptotics satisfy,
@@
$\qquad\qquad+C_{2}+f_{2}(\alpha_{F_{p}})+O\big{(}s^{\min\{{\log(\beta/|\beta_{2}|)}/{\log\beta},1-\varepsilon\}})$
$=\frac{1}{2\log\beta}(\log n)^{2}\left(1+O\left(\frac{1}{\log n}\right)\right)$
$\qquad\qquad+\left(\log n+\log\log\beta+O\left(\frac{1}{\log n}\right)\right)\left(\frac{2\gamma}{\log\beta}-\frac{\log\lambda}{\log\beta}-\frac{1}{2}\right)$
$\qquad\qquad\qquad+C_{2}+\psi_{4}\left(\frac{\log n}{\log\beta}\right)+O\left(n^{-\min\{{\log(\beta/|\beta_{2}|)}/{\log\beta},1-\varepsilon\}}\right)$
$=\frac{(\log n)^{2}}{2\log\beta}+(\log n)\left(\left(\frac{2\gamma}{\log\beta}-\frac{\log\lambda}{\log\beta}-\frac{1}{2}\right)\right)$
$\qquad\qquad+\left(\log\log\beta\right)\left(\frac{2\gamma}{\log\beta}-\frac{\log\lambda}{\log\beta}-\frac{1}{2}\right)$
$\qquad\qquad\qquad+C_{2}+\psi_{4}\left(\frac{\log n}{\log\beta}\right)+O\left(\frac{1}{\log n}\right),$
where $\psi_{4}(x)$ is an explicitly determinable $1$-periodic function. Thus, we have
$F_{P}(e^{-\alpha_{F_{P}}})=\psi_{5}\left(\frac{\log n}{\log\beta}\right)n^{\left(\frac{\log n}{2\log\beta}+\frac{2\gamma}{\log\beta}-\frac{\log\lambda}{\log\beta}-\frac{1}{2}+O\left(\frac{1}{\log n}\right)\right)}\left(1+O\left(\frac{1}{\log n}\right)\right),$
where $\psi_{5}(x)$ is the $1$-periodic function satisfying
$\psi_{5}\left(\frac{\log n}{\log\beta}\right):=\exp\left(\left(\log\log\beta\right)\left(\frac{2\gamma}{\log\beta}-\frac{\log\lambda}{\log\beta}-\frac{1}{2}\right)+C_{2}+\psi_{4}\left(\frac{\log n}{\log\beta}\right)\right).$
Combing the asymptotic for $F_{P}(e^{-\alpha_{F_{P}}})$ with equations (20) and (21) gives the desired result for $p_{P}(n)$, with $\psi_{p_{P}}(x):={e^{\frac{1}{\log\beta}}\psi_{1}(x)}/{\sqrt{2\pi\log\varphi}}.$
@@ Acknowledgements
This work was started during a short visit of M. Coons to Aarhus University; he thanks their faculty and staff for their hospitality and support. S. Kristensen and M. L. Laursen are supported by the Independent Research Fund Denmark (Grant ref. 1026-00081B) and the Aarhus University Research Research Foundation (Grant ref. AUFF-E-2021-9-20). The idea for this work came about when M. Coons was a master’s student working under Klaus Kirsten, and certainly this work benefitted from several conversations with Klaus and also Markus Hunziker, which occurred nearly twenty years ago! M. Coons gives a very belated thanks to both of them.

### v3

@@ 4. Partitions over general linear recurrences with a dominant root
We adopt the terminology of Proposition 8 for the rest of this section along with the definitions of the real numbers $\lambda=\lambda_{1},\lambda_{2},\ldots,\lambda_{r}$, which satisfy
@@
$\zeta_{P}(z)=\frac{1}{\log\beta}\cdot\frac{1}{z}-\frac{\log\lambda}{\log\beta}-\frac{1}{2}+\left(\frac{(\log\lambda)^{2}}{2\log\beta}+\frac{\log\lambda}{2}+\frac{\log\beta}{12}+C_{1}\right)z+O(z^{2}).\qed$
@@
$s^{-z}\Gamma(z)\zeta(1+z)=\frac{1}{z^{2}}+(2\gamma-$ $\log s)\frac{1}{z}$
$+\left(\frac{1}{2}\left(\gamma^{2}+\frac{\pi^{2}}{6}\right)+\gamma^{2}-\gamma_{1}-2\gamma\log s+\frac{(\log s)^{2}}{2}\right)+O(z),$
@@
$\qquad-\left(\frac{\log\lambda}{\log\beta}+\frac{1}{2}\right)(2\gamma-\log s)$
$\qquad\qquad+\left(\frac{(\log\lambda)^{2}}{2\log\beta}+\frac{\log\lambda}{2}+\frac{\log\beta}{12}+C_{1}\right)$
$=\frac{(\log s)^{2}}{2\log\beta}+\left(\frac{\log\lambda}{\log\beta}+\frac{1}{2}-\frac{2\gamma}{\log\beta}\right)\log s+C_{2}.$
The contributions coming from the rest of the poles are as discussed before the statement of Lemma 4. ∎
On inspection, one notices that the dominant asymptotic terms of $\log F_{P}(e^{-s})$ are precisely the dominant asymptotic terms of $\log F_{2}(e^{-s})$, the function related to the Fibonacci partitions, after substituting $\beta$ for $\varphi$. Of course, this is not so surprising, as the Fibonacci numbers $F_{k}$ are just a special case of the more general sequence $P_{k}$. The property of note, here, is regarding the associated saddle point. Since the leading order behaviour is the same, the saddle points satisfy the same leading order asymptotics. In particular, Proposition 6 and its corollaries hold for the saddle point $\alpha_{F_{P}}$ related to non-distinct partitions of $n$ over $P_{k}$, and so also then, do (23) and (24). We use these results, as well as this notation, below.
@@
By Proposition 9, Proposition 6 and its corollaries, the small-$\alpha_{F_{P}}$ asymptotics, or, equivalently, the large-$n$ asymptotics satisfy,
@@
$\qquad\qquad+C_{2}+f_{2}(\alpha_{F_{p}})+O\Big{(}\alpha_{F_{p}}^{\min\{{\log(\beta/|\beta_{2}|)}/{\log\beta},1-\varepsilon\}}\Big{)}$
$=\frac{\log n}{2\log\beta}\left(\log n-2\log\log n-\frac{4\psi_{3}\left(\frac{\log n}{\log\varphi}\right)}{\log\beta}+2\log\log\beta+4\gamma-2\log\lambda-\log\beta\right)$
$\qquad+\frac{\log\log n}{2\log\beta}\left(\log\log n+2-\frac{4\psi_{3}\left(\frac{\log n}{\log\beta}\right)}{\log\beta}+2\log\log\beta-4\gamma+2\log\lambda+\log\beta\right)$
$\qquad\qquad+\frac{(\log\log\beta)^{2}}{2\log\beta}-\frac{\psi_{3}\left(\frac{\log n}{\log\beta}\right)}{(\log\beta)^{2}}+(1+2\gamma-\log\lambda)\frac{\log\log\beta}{\log\beta}-\frac{1}{2}\log\log\beta$
$\qquad\qquad\qquad+C_{2}+\psi_{4}\left(\frac{\log n}{\log\beta}\right)+O\left(\frac{(\log\log n)^{2}}{\log n}\right),$
where $\psi_{3}(x)$ and $\psi_{4}(x)$ are explicitly computable $1$-periodic functions that are analogous to $\psi_{0}(x)$ and $\psi_{1}(x)$, respectively. Thus, we have
$F_{P}(e^{-\alpha_{F_{P}}})=\psi_{5}\left(\frac{\log n}{\log\beta}\right)n^{c(n)}(\log n)^{d(n)}\left(1+O\left(\frac{(\log\log n)^{2}}{\log n}\right)\right),$
where $\psi_{5}(n)$ is the $1$-periodic positive function
$\psi_{5}(x)$ $:=\exp\Bigg{(}\frac{(\log\log\beta)^{2}}{2\log\beta}-\frac{\psi_{3}(x)}{(\log\beta)^{2}}+(1+2\gamma-\log\lambda)\frac{\log\log\beta}{\log\beta}-\frac{1}{2}\log\log\beta+C_{2}+\psi_{4}(x)\Bigg{)},$
$c(n)$ $:=\frac{1}{2\log\beta}\left(\log n-2\log\log n-\frac{4\psi_{3}\left(\frac{\log n}{\log\varphi}\right)}{\log\beta}+2\log\log\beta+4\gamma-2\log\lambda-\log\beta\right),$
$d(n)$ $:=\frac{1}{2\log\beta}\left(\log\log n+2-\frac{4\psi_{3}\left(\frac{\log n}{\log\beta}\right)}{\log\beta}+2\log\log\beta-4\gamma+2\log\lambda+\log\beta\right).$
Combing the asymptotic for $F_{P}(e^{-\alpha_{F_{P}}})$ with equations (23) and (24) gives the desired result. ∎
@@ Acknowledgements
This work was started during a short visit of M. Coons to Aarhus University; he thanks their faculty and staff for their hospitality and support. S. Kristensen and M. L. Laursen are supported by the Independent Research Fund Denmark (Grant ref. 1026-00081B) and the Aarhus University Research Foundation (Grant ref. AUFF-E-2021-9-20), and M. Coons is supported by the David W. and Helen E. F. Lantis Endowment. The idea for this work came about when M. Coons was a master’s student working under Klaus Kirsten—this work benefitted from several conversations with Klaus Kirsten and also Markus Hunziker, which occurred nearly twenty years ago! M. Coons gives a hearty, and very belated, thanks to both of them.